OpenStudy (anonymous):

Solve the following equation for a. Please show all of your work c=2a-15/b

OpenStudy (anonymous):

Whenever you're asked to 'solve' for a variable, all that's wanted is that you perform the required operations so that the variable is on one side (alone) with everything else on the other. Here, 'a' is the subject, so I would first add 15/b to both sides,\[c+\frac{15}{b}=2a\]and then divide both sides by 2 to isolate 'a',\[\frac{c+\frac{15}{b}}{2}=a\]You can then simplify the left-hand side of the last equation,\[\frac{c+\frac{15}{b}}{2}=\frac{c}{2}+\frac{15/b}{2}=\frac{c}{2}+\frac{15}{b} \times \frac{1}{2}=\frac{c}{2}+\frac{15}{2b}\]so that\[a=\frac{c}{2}+\frac{15}{2b}\]

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thanx